# Supervised Learning-enhanced Multi-Group Actor Critic for Live-stream Recommendation

### v1

@@ 3.3.2. Critic Network
In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of the critic networks. Therefore, we seamlessly introduce supervised learning (e.g. multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure. Specifically, we split the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
@@
Since the distribution of live-stream (short video) watch time changes drastically over time, the time gain reward $r$ in Eq. 2 becomes difficult to learn. In this work, we propose a distribution discretization method to improve reward learning.
Formally, we divide the live-stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live-stream (short video) watch time. Let $y$ denote the actual live-stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$. The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$. Then, we have a linear reconstruction layer to reconstruct $y$ from $\delta$:
@@
Let $o(t)_{l}\in\mathbb{R}^{N_{l}+1}$, $o(t)_{v}\in\mathbb{R}^{N_{v}+1}$ be one-hot vectors representing the time bins in which the real live-stream (short video) watch time of sample $(s_{t},a_{t},r_{t})$ falls. Note that a separate bin is set for the case $a_{t}=0$. Then, the RPN $R_{\theta_{i}}$ can be modeled by multi-task neural networks:
@@
where $y_{l}^{st},y_{l}^{end}\in\mathbb{R}^{N_{l}+1}$ predefined left (right) boundary vectors for the live-stream watch time bin, $y_{v}^{st},y_{v}^{end}\in\mathbb{R}^{N_{v}+1}$ are predefined left (right) boundary vectors for the short video watch time bin. Note that introducing the posterior one-hot vectors $o(t)_{l}$ and $o(t)_{v}$ in Eq. 8 will not affect the calculation of $Q^{\prime}_{\phi_{i}}(s_{t+1},\cdot)$, because we have already reserved $r_{t+1}$ in the dataset $\mathcal{D}$. Therefore, $o(t+1)_{l}$ and $o(t+1)_{v}$ can be easily obtained from $r_{t+1}$ to compute $R^{\prime}_{\theta_{i}}(s_{t+1},\cdot)$.
As shown in Eq. 8, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin. We then reconstruct the predicted live-stream and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively. Finally, we obtain the predicted normalized reward, $R_{\theta_{i}}(s_{t},a_{t})$. To optimize the predicted watch time proportions, we use Huber Loss as follows:
@@
where $\delta_{l}$, $\delta_{v}$ are the actual proportion labels for live-stream and short video watch times, respectively.
The reason for using Huber loss on the time proportion $\delta$ instead of the actual watch time $y$ is that the time proportion lies within the range $[0,1]$, which results in smaller gradients in the neural network. Additionally, the variance of $\delta$ is much smaller than that of the original watch time $y$. Consequently, the outputs of the learned reward networks ($f_{\Gamma_{i}}$ and $g_{\Theta_{i}}$) will exhibit smaller variances. Hence, reward learning with the discretization of watch time distributions can be viewed as a novel variance reduction technique.
@@ 3.3.3. Actor Network
Moreover, we observe that large values of $\hat{Q}(s_{t},a_{t})$ may cause instability in actor training, which can ultimately lead to policy deterioration, where $\forall t,\;\pi(a_{t}=1|s_{t})>\pi(a_{t}=0|s_{t})$ or $\pi(a_{t}=1|s_{t})<\pi(a_{t}=0|s_{t})$. To address this, we apply softmax normalization to $\hat{Q}(s_{t},a_{t})$ to obtain $\hat{Q}_{norm}=softmax(\hat{Q}(s_{t},a_{t}))$, and propose a modified actor loss function:
@@
We note that the actor loss in Eq. LABEL:eq:15 is a standard cross-entropy loss function. From the perspective of knowledge distillation (Hinton, 2015), the actor distills policy knowledge from more complex critic networks. The teacher critic networks guide the more lightweight student actor network to adjust and converge to an optimal policy. In practice, we only need to deploy the actor network in the online live-stream recommendation system, which significantly reduces computational complexity and improves real-time respond speed.
@@ 3.4. Online Exploration and Deployment
We deploy the SL-MGAC algorithm in our recommendation system, and the overall system architecture is shown in Fig. 4. The online RL agent collects real-time user interaction logs, while the offline model trainer optimizes the SL-MGAC model in an off-policy manner using streaming training data. Moreover, the offline trainer sends the latest model parameters to update the online deployed actor network in real-time.
@@ 4. Experiments
We conduct both offline evaluation on a real-world dataset collected from our recommendation system and online A/B test experiments with SL-MGAC and the baselines.
@@ 4.1. Dataset
Due to the lack of publicly available datasets for recommendation decisions involving live-streaming injections in short video feeds, we collect a dataset from our recommendation system with 10,000 users through random sampling over a full day, ensuring that the proportions of different user groups are similar to those in real online data. In total, we have over 180,000 samples, which are split into training and test sets with a 4:1 ratio.
@@ 4.2.1. Baselines
  • DQN (Mnih et al., 2015). A deep neural network algorithm for Q-learning that introduces the technique of target network updates.
@@
  • TD3 (Fujimoto et al., 2018). A modified version of DDPG (Lillicrap, 2015) that addresses function approximation errors using three techniques: Clipped Double Q-Learning, Delayed Policy Updates, and Target Policy Smoothing.
@@
Note that in our optimal live-stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique into TD3 and TD3-BC.
@@ 4.2.2. Variations of our model
  • SL-MGAC (w/o Q-norm): An SL-MGAC variant without Q value normalization in the actor loss.
@@
  • SL-MGAC-vanilla: An SL-MGAC variant with the vanilla Q label in the critic loss, as shown below, where $p(s_{t+1},\cdot)$ is the action probability of $s_{t+1}$ output by the actor network.
@@ 4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with an embedding layer of size 5000. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods. The detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of Appendix B.
@@ 4.4. Offline Policy Evaluation
We follow the approach in (Xue et al., 2023) and adopt a commonly used offline policy evaluation method, namely Normalized Capped Importance Sampling (NCIS) (Swaminathan and Joachims, 2015), to evaluate performance. The evaluation metric is the cumulative reward across all trajectories of test users. Note that the reward for each sample is defined in Eq. 3.
@@
The offline policy evaluation results are shown in Table 1. Compared to the supervised learning method L2R, most RL methods achieve higher cumulative rewards, except for DQN. SL-MGAC significantly outperforms all other methods, demonstrating that SL-MGAC can achieve higher long-term rewards in the complex live-stream & short video mixed recommendation system with drastically changing data distributions. Furthermore, SL-MGAC outperforms SL-MGAC-vanilla, suggesting that incorporating the actor network in the Q label calculation of Eq. 15 may interfere with critic learning to some extent, slightly affecting the performance of SL-MGAC-vanilla. Additionally, SL-MGAC converges faster than SL-MGAC-sep during training, as shown in Fig. 8 in Appendix C, and achieves a higher cumulative reward in offline evaluation. This may be due to the information loss caused by detaching $\hat{Q}_{norm}$, which can lead to slower convergence for SL-MGAC-sep.
@@ 4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of live-stream daily active users (DAU) and watch time relative to the baseline. The baseline uses the SAC framework for online live-stream injection. The online A/B test results are shown in Table 3. SL-MGAC achieves the largest improvement in live-stream watch time while also increasing live-stream DAU. Although SL-MGAC-0 shows a larger improvement in live-stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live-streams, which may negatively impact the long-term user experience for most users. Therefore, this demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live-streams in response to user requests.
Methods Live-stream DAU Live-stream Watch Time
L2R +0.321% -0.558%
SL-MGAC-0 +3.928% -2.934%
SL-MGAC +2.616% +7.431%
Table 3. Online A/B Test Performance of compared methods.
Since the scale of live-stream supply and user interaction behaviors change drastically over time, more streamers tend to broadcast live during peak hours. As shown in Fig. 9 in Appendix C, SL-MGAC can adaptively adjust the live-stream injection strategy to increase live-stream exposure during peak hours, while reducing exposure during low-traffic periods.
@@ 4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 7, the live-stream injection ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live-stream injection ratio refers to the proportion of requests with action $a=1$.
Figure 7. The trend curve of online live-stream injection ratio during a whole day.
We calculate the amplitudes of the live-stream injection ratio within sliding time windows of 20 minutes. The amplitude density is shown in Fig. 10 in Appendix C. The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers. In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, which demonstrates the superior online model stability of SL-MGAC.
@@ 5.1. RL in Recommendation Systems
Reinforcement learning (RL) aims to optimize long-term cumulative rewards over time, which has attracted significant attention in recommendation systems research in recent years (Afsar et al., 2022). Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend entire item lists, where the number of candidate items can be large. BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation. The work in (Miao et al., 2021) explores the use of off-policy RL for multi-session personalized search ranking. CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving user experience. Additionally, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
@@ 5.2. RL divergence and instability
Recently, there has been growing literature focusing on RL divergence and instability. (Van Hasselt et al., 2016) introduces Double DQN to mitigate the maximization bias problem in DQN. (Fujimoto et al., 2018) proposes the Clipped Double Q-Learning technique in TD3 to reduce the overestimation of Q-values. (Mao et al., 2018) introduces a bias-free, input-dependent baseline for the policy gradient algorithm (Sutton et al., 1999) to reduce variance and improve training stability. (Dohare et al., 2023) investigates policy collapse in a 2-state MDP and finds that L2 regularization and the non-stationary Adam optimizer (Kingma, 2014) are both effective in alleviating RL instability. (Yue et al., 2024) theoretically analyzes the causes of RL divergence and applies Layer Normalization to mitigate RL divergence and instability.

### v2

@@ 3.3.2. Critic Network
In practice, $Q_{label}(s_{t+1})$ may be dominated by $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ during critic learning, which can affect the performance of critic networks. Specifically, due to the instability and divergence issues, target networks often predict inaccurate $\hat{Q}^{\prime}(s_{t+1},a_{t+1})$ values, which is much larger than the $r(s_{t},a_{t})$ term. Then the $r(s_{t},a_{t})$ term could not provide any information to guide the critic learning. Therefore, we seamlessly introduce supervised learning (e.g. multi-task learning (Zhang and Yang, 2021)) into the critic learning procedure. Specifically, we divide the critic network $Q_{\phi_{i}}$ (or $Q^{\prime}_{\phi_{i}}$) into two components: a Reward Prediction Network (RPN) and a Q Residual Network (QRN), as follows:
@@
Since the distribution of live stream (short video) watch time changes drastically over time, the time gain reward $r$ in Eq. 4 becomes difficult to learn. In this work, we propose a distribution discretization method to improve reward learning.
Formally, we divide the live stream and short video watch time distribution into $N_{l}$ and $N_{v}$ non-overlapping bins, where each bin represents an interval of live stream (short video) watch time. Let $y$ denote the actual live stream or short video watch time, which falls into the time bin $[y_{st},y_{end}]$. The proportion of $y$ within this bin is $\delta=\frac{y-y_{st}}{y_{end}-y_{st}}\in[0,1]$. Then, we have a linear reconstruction layer to reconstruct $y$ from $\delta$:
@@
Let $o(t)_{l}\in\mathbb{R}^{N_{l}+1}$, $o(t)_{v}\in\mathbb{R}^{N_{v}+1}$ be one-hot vectors representing the time bins in which the real live stream (short video) watch time of sample $(s_{t},a_{t},r_{t})$ falls. Note that a separate bin is set for the case $a_{t}=0$. Then, the RPN $R_{\theta_{i}}$ can be modeled by multi-task neural networks:
@@
where $y_{l}^{st},y_{l}^{end}\in\mathbb{R}^{N_{l}+1}$ predefined left (right) boundary vectors for the live stream watch time bin, $y_{v}^{st},y_{v}^{end}\in\mathbb{R}^{N_{v}+1}$ are predefined left (right) boundary vectors for the short video watch time bin. Note that we introduce the posterior one-hot vectors $o(t)_{l}$ and $o(t)_{v}$ in Eq. 9 will not affect the calculation of $Q^{\prime}_{\phi_{i}}(s_{t+1},\cdot)$, because we have already reserved $r_{t+1}$ in the dataset $\mathcal{D}$. Therefore, $o(t+1)_{l}$ and $o(t+1)_{v}$ can be easily obtained from $r_{t+1}$ to compute $R^{\prime}_{\theta_{i}}(s_{t+1},\cdot)$.
As shown in Eq. 9, we employ two separate MLPs, $f_{\Gamma_{i}}(s_{t},a_{t})$ and $g_{\Theta_{i}}(s_{t},a_{t})$, to predict the proportions within a time bin. We then reconstruct the predicted live streaming and short video watch times, $F_{\Gamma_{i}}(s_{t},a_{t})$ and $G_{\Theta_{i}}(s_{t},a_{t})$, respectively. Finally, we obtain the predicted normalized reward, $R_{\theta_{i}}(s_{t},a_{t})$. To optimize the predicted watch-time proportions, we use Huber loss as follows:
@@
where $\delta_{l}$, $\delta_{v}$ are the actual proportion labels for live stream and short video watch times, respectively.
The reason for using the Huber loss on the time ratio $\delta$ instead of the actual watch time $y$ is that the time ratio is within the range $[0,1]$, resulting in smaller gradients in the neural network. Furthermore, the variance of $\delta$ is much smaller than that of the original watch time $y$. Consequently, the outputs of the learned reward networks ($f_{\Gamma_{i}}$ and $g_{\Theta_{i}}$) will exhibit smaller variances. Hence, reward learning with the discretization of watch time distributions can be viewed as a novel variance reduction technique.
@@ 3.3.3. Actor Network
Moreover, we observe that high values of $\hat{Q}(s_{t},a_{t})$ can cause instability in actor training, which can ultimately lead to policy deterioration, where $\forall t,\;\pi(a_{t}=1|s_{t})>\pi(a_{t}=0|s_{t})$ or $\pi(a_{t}=1|s_{t})<\pi(a_{t}=0|s_{t})$. To address this, we apply softmax normalization to $\hat{Q}(s_{t},a_{t})$ to obtain $\hat{Q}_{norm}=softmax(\hat{Q}(s_{t},a_{t}))$, and propose a modified actor loss function:
@@
We note that the actor loss in Eq. LABEL:eq:15 is a standard cross-entropy loss function. From the perspective of knowledge distillation (Hinton, 2015), the actor distills policy knowledge from more complex critic networks. The teacher critic networks guide the more lightweight student actor network to adjust and converge to an optimal policy. In practice, we only need to deploy the actor network in the online live stream recommendation system, which significantly reduces computational complexity and improves real-time response speed.
@@ 3.4. Online Exploration and Deployment
We implemented the SL-MGAC algorithm in our recommendation system, and the overall architecture of the system is shown in Fig. 3. The online RL agent collects real-time user interaction logs, while the offline model trainer optimizes the SL-MGAC model in an off-policy manner using streaming training data. Moreover, the offline trainer sends the latest model parameters to update the online deployed actor network in real-time.
@@ 4. Experiments
We perform both offline evaluation on a real-world dataset collected from our recommendation system and online A/B test experiments with SL-MGAC and the baselines.
@@ 4.1. Dataset
Due to the lack of publicly available datasets for recommendation decisions involving live streaming injections in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data. In total, we have over 1,800,000 samples, which are divided into training and test sets with a 4:1 ratio. In addition, the online environment contains more than 100 million app users for AB-Tests.
@@ 4.2.1. Baselines
  • DQN (Mnih et al., 2015). A deep neural network algorithm for Q-learning that introduces the technique of updating the target network.
@@
  • TD3 (Fujimoto et al., 2018). A modified version of DDPG (Lillicrap, 2015) addresses function approximation errors using three techniques: Clipped Double Q-Learning, Delayed Policy Updates, and Target Policy Smoothing.
@@
Note that in our optimal live stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
@@ 4.2.2. Variations of our model
  • SL-MGAC (w/o Q-norm): An SL-MGAC variant without Q-value normalization in actor loss.
@@
  • SL-MGAC-vanilla: An SL-MGAC variant with the vanilla Q label in the loss of the critic, as shown below, where $p(s_{t+1},\cdot)$ is the probability of action of the output of $s_{t+1}$ by the actor network.
@@ 4.3. Implementation Details
To ensure fairness among all compared methods, we use a consistent feature extraction module with a 5000-size embedding layer. In addition to the embedding layer, the feature extraction module includes a 2-layer MLP with hidden sizes of [256, 128]. The batch size and number of epochs for all methods are 2048 and 500, respectively. The learning rate for the embedding layer is set to 1e-5, while the learning rate for other hidden parameters is 1e-3. The discount factor $\gamma$ is set to 0.9 for all methods. Detailed network architecture and parameter settings for SL-MGAC are provided in Table 4 of the Appendix A.
@@ 4.4. Offline Policy Evaluation
We follow the approach in (Xue et al., 2023) and adopt a commonly used offline policy evaluation method, namely Normalized Capped Importance Sampling (NCIS) (Swaminathan and Joachims, 2015), to evaluate performance. The evaluation metric is the cumulative reward across all trajectories of test users.
@@
The results of the offline policy evaluation are shown in Table 1. Compared to the supervised learning method L2R, most RL methods achieve higher cumulative rewards, except for DQN. SL-MGAC significantly outperforms all other methods, demonstrating that SL-MGAC can achieve higher long-term rewards in the complex live stream & short video mixed recommendation system with drastically changing data distributions. Furthermore, SL-MGAC outperforms SL-MGAC-vanilla, suggesting that incorporating the actor network in the Q label calculation of Eq. 16 may interfere with the learning of the critic to some extent, slightly affecting the performance of the SL-MGAC-vanilla.
@@ 4.6. Parameter Sensitivity
Methods live stream DAU live stream Watch Time Video Watch Time App Usage Duration User Retention
L2R +0.321% -0.558% -0.238% -0.178% -0.154%
Dummy +51.3% +34.6% -2.216% -1.062% -0.439%
SL-MGAC-0 +3.928% -2.934% -0.153% -0.117% -0.102%
SL-MGAC +2.616% +7.431% +0.197% +0.121% +0.086%
Table 3. Online A/B Test Performance of compared methods.
@@ 4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of live stream daily active users (DAU), live stream watch time, video watch time, app usage duration and user retention relative to the baseline. The baseline uses the SAC framework for live stream injection. The results of the online A / B test are shown in Table 3.
Note that the Dummy method injects a live stream for each request. We can see that SL-MGAC achieves the largest improvement in live stream watch time while also increasing the app usage duration and user retention than other methods. It demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live streams in response to user requests. Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may negatively impact the long-term user experience for most users. Moreover, the dummy method can greatly improve live streaming DAU and watch time, but significantly affect app usage time and user retention.
@@ 4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 6, the live stream injection ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live stream injection ratio refers to the proportion of requests with action $a=1$.
Figure 6. The trend curve of online live stream injection ratio during a whole day.
We calculate the amplitudes of the live stream injection ratio within sliding time windows of 20 minutes. The amplitude density is shown in Fig. 8 in Appendix B. The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers. In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, demonstrating the superior online model stability of SL-MGAC.
5. Discussion
The intrinsic instability and divergence of RL limits its application in real-world industrial recommendation systems. The proposed SL-MGAC method improves the robustness and stability of the RL model and achieves significant offline and online performance in the live stream and short video mixed recommendation scenario. However, SL-MGAC only learns from data from adjacent states $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1},r_{t+1})$, which limits its ability to learn information in much longer sequences, such as the daily-level user preference, rather than session-level user preference, since the capture of daily user preference information is very helpful for improving user engagement and user retention. Learning live stream policy from a long sequence of user behaviors would lead to our future work.
Furthermore, we emphasize that more sophisticated methods can also be employed to divide users into multiple groups and provide prior information about user preference, such as traditional clustering methods, e.g. k-means, neural clustering of VQ-VAE (Van Den Oord et al., 2017), online clustering of SwAV (Caron et al., 2020) by the Sinkhorn-Knopp algorithm (Cuturi, 2013), etc. We find that the simple MG-SD module achieves significant performance and fast real-time response spped in practice; we will lead these complicated group partitioning methods in our further research.
@@ 6.1. RL in Recommendation Systems
Reinforcement learning (RL) aims to optimize cumulative long-term rewards over time, which has attracted significant attention in the research of recommendation systems in recent years (Afsar et al., 2022). Methods such as SLATEQ (Ie et al., 2019), GeMS (Deffayet et al., 2023), and HAC (Liu et al., 2023) use RL to recommend complete item lists, where the number of candidate items can be large. BatchRL-MTF (Zhang et al., 2022), RLUR (Cai et al., 2023), and UNEX-RL (Zhang et al., 2024) leverage RL to model the multi-rank score aggregation process, optimizing the weights for score aggregation. The work in (Miao et al., 2021) explores the use of off-policy RL for personalized search ranking in multiple sessions. CrossDQN (Liao et al., 2022) introduces a RL-based approach for ad allocation in a feed, aiming to maximize revenue while improving the user experience. Furthermore, traditional RL methods such as DQN (Mnih, 2013; Mnih et al., 2015), Double DQN (Van Hasselt et al., 2016), SAC (Haarnoja et al., 2018), DDPG (Lillicrap, 2015), and TD3 (Fujimoto et al., 2018) serve as backbones in real-world RL applications for recommendation systems.
@@ 6.2. RL divergence and instability
Recently, there has been a growing literature that focuses on RL divergence and instability. (Van Hasselt et al., 2016) introduces Double DQN to mitigate the maximization bias problem in DQN. (Fujimoto et al., 2018) proposes the Clipped Double Q-Learning technique in TD3 to reduce the overestimation of Q values. (Mao et al., 2018) introduces a bias-free, input-dependent baseline for the policy gradient algorithm (Sutton et al., 1999) to reduce variance and improve the stability of training. (Dohare et al., 2023) investigates policy collapse in a 2-state MDP and finds that L2 regularization and the non-stationary Adam optimizer (Kingma, 2014) are both effective in alleviating RL instability. (Yue et al., 2024) theoretically analyzes the causes of RL divergence and applies Layer Normalization to mitigate RL divergence and instability.